# Learning Shared Representations from Unpaired Data

### v1

@@ 5 Experiments
The remainder of this section is organized as follows. Sec. 5.1 evaluates retrieval performance. Sec. 5.2 presents a range of applications of universality. To better understand the contribution of each component in SUE, we investigate the effects of various components. In Sec. 5.3, we examine the impact of incorporating additional paired and unpaired samples, revealing that unpaired data holds greater potential than previously recognized. Sec. 5.4 presents an ablation study assessing the roles of SUE’s components (SE, CCA, and MMD), demonstrating that the novel integration of these techniques is central to SUE’s effectiveness. Due to space constraints, we refer to App. A for additional results. Further details on the experimental setup, implementation, and training procedures are provided in Appendix D.
@@
To evaluate SUE’s performance, we use several paired datasets. To provide informative qualitative results and facilitate an intuitive understanding of the universal embedding concept, we use three vision-language datasets ([Uncaptioned image]): Flickr30k Plummer et al. (2015), MSCOCO Lin et al. (2014), and Polyvore Han et al. (2017); as well as a vision-vision dataset ([Uncaptioned image]): Edges2Shoes Isola et al. (2017). For the generation task, we use another vision-language dataset ([Uncaptioned image]): caption-FFHQ Kazemi and Sullivan (2014). See App. D.1 for further details on the datasets and preprocessing (e.g., the unimodal models used and adaptation to weakly-paired settings).
@@
To the best of our knowledge, this is one of the very first works to exploit weakly-paired data (i.e., very few pairs). As such, no direct baselines exist for this setting. To contextualize results, we compare SUE with the contrastive method designed for fully paired data Radford et al. (2021); Xu et al. (2021); Elizalde et al. (2023), training an MLP on pre-trained unimodal features using the same number of pairs. We also consider domain confusion (App. A.1), which performs notably worse in multimodal scenarios.
Table 1: Retrieval results. Results with few paired samples on vision-language ([Uncaptioned image]) and vision-vision ([Uncaptioned image]) datasets from each modality to another: image-to-text (I2T), text-to-image (T2I), edges-to-shoes (E2S), shoes-to-edges (S2E); by SUE and Contrastive. The Imp. column states the relative mean improvement of SUE over Contrastive learning. Using the same small number of pairs, SUE significantly outperforms the popular paired method. SUE substantially relies on unpaired data.
#paired SUE (ours) Contrastive Imp.
R@1 R@5 R@10 R@1 R@5 R@10
MSCOCO 100 I2T 5.75 21.50 34.25 1.50 8.50 13.00 +257.20%
[Uncaptioned image] T2I 5.25 18.25 33.25 0.80 5.80 12.20
Flickr30k 500 I2T 4.25 19.75 32.00 3.00 9.50 16.20 +103.32%
[Uncaptioned image] T2I 5.75 22.00 32.75 2.50 9.80 15.00
Polyvore 500 I2T 6.00 22.75 32.25 3.20 13.8 22.5 +55.67%
[Uncaptioned image] T2I 4.75 20.75 32.00 4.00 11.50 23.00
Edges2Shoes 50 E2S 4.00 16.00 25.25 1.0 5.50 14.00 +200.51%
[Uncaptioned image] S2E 3.50 17.00 27.00 0.80 6.00 12.80
@@ 5.1 Retrieval
Tab. 1 reports retrieval scores for SUE and contrastive learning on several datasets using a randomly sampled test set of 400 examples. Following prior work Radford et al. (2021); Elizalde et al. (2023), we use the Recall@$k$ metric. The results demonstrate SUE’s ability to capture cross-modal semantics with as few as 100 pairs for MSCOCO, 500 for Flickr30k and Polyvore, and just 50 for Edges2Shoes. Remarkably, these results are also comparable with the ones reported by Huang et al. (2022) on the Flickr30k T2I task [^1] R@1 10.3; R@5 25.1; R@10 34.0., despite MACK using a paired segmentation-to-text model (see App. A). In App. A.4, we further analyze how different unimodal encoders affect SUE’s retrieval performance.
@@
In Fig. 4a we supplement the quantitative results with several qualitative examples. Additional qualitative results can be found in App. A.7. These results illustrate how SUE successfully captures semantic relationships across modalities, even in cases where the exact pair is not among the first neighbors (leading to a zero Recall@$k$ score).
In Fig. 2, we consider a custom caption scenario, where we retrieve images of a more general query sentence that cannot be found in the dataset. These examples further support SUE’s ability to capture these cross-modal semantics. These qualitative examples also depict the challenge of reliably measuring semantic coherence across modalities in retrieval tasks using the recall@$k$ metric. Therefore, in App. C we consider a new soft variant of Recall@$k$, which follows previous works Hessel et al. (2021) and uses CLIP score as a measure of semantic alignment between image-text pairs.
@@ 5.2 Applications of Universality
Fig. 4b depicts several examples of text-to-image generation with minimal text-image correspondence. In particular, we use a pre-trained GAN inversion model Wang et al. (2022) trained on the FFHQ dataset Kazemi and Sullivan (2014) and train a converter model that maps from the universal image embeddings to the GAN’s latent space. During inference, we pass universal text embeddings to the converter and generate face images from its output using the GAN decoder. Noteworthy, both the GAN and the converter were trained exclusively on images, with few text-image pairs available only during SUE’s training. For more details on the training process of the converter model, please refer to App. D.3.
@@
Fig. 4c presents the results of emergent arithmetic operations on the universal embedding. Notably, these examples require no additional training - we simply perform vector summation of text and image embeddings and generate the image using the pre-trained converter and GAN. This approach aligns with techniques demonstrated using CLIP Ramesh et al. (2022) but with limited text-image pairing. This summation in the vector space translates into an intuitive integration of semantics, making it, in a sense, a semantic vector space, while almost no text-image pairs were available during the construction of the shared embedding. This showcases the SUE’s cross-modality capturing abilities, even in the absence of large paired datasets.
@@
In App A.2 we further demonstrate SUE’s ability to capture high-level cross-modal, using a zero-shot scenario - acting as a classifier without being explicitly trained for classification.
@@
In App A.3 we demonstrate the effectiveness of universality for the classification of an unlabeled domain with minimal correspondence to a labeled domain.
@@ 5.4 Ablation Study
Tab. 2 presents the results of various ablations of SUE, from the original unimodal representations to the complete SUE pipeline. Alongside these empirical results, we provide UMAP visualizations of SUE’s three steps in Fig. 13 in App. A.9. The main outcome from these results is that SE has a pivotal role in revealing the universal embedding. Attempts to align the original unimodal representations yield mostly trivial retrieval results (see left column of each dataset in Tab. 2). However, applying the same alignment approach (i.e., CCA alongside MMD) on the SE space, significantly improves the retrieval performance. This underscores SE’s essential role in extracting the modality-invariant representations. This is further supported by the results in Tab. 9 in App. A.8, where replacing the SE step with an Auto-Encoder significantly decreases the performance. These results support the existence of the universal embedding. SUE successfully learns a meaningful cross-modal semantic representation with only a minimal number of paired samples, used solely to facilitate linear cross-modal alignment. See extended ablation study in App. A.8.
@@ 6 Conclusion and Future Work
In doing so, we challenge the prevailing approach in multimodal learning, which relies heavily on extremely large amounts of paired samples. SUE, operating with weakly-paired data, in the sense of very few pairs, yields meaningful results across various applications, especially highlighted by the comparison to contrastive learning. Many additional tasks and modalities remain to be explored, and we consider these an intriguing avenue for future work. Although these results do not yet surpass current state-of-the-art models trained with a huge amount of pairs, they provide compelling evidence for the hidden potential of unpaired data, demonstrating the ability to learn a shared representation space primarily from unpaired examples.
Importantly, our work does not yet aim to replace existing multimodal models trained with large sets of paired data. Rather, it introduces a new concept for settings where such supervision is unavailable. Interestingly, the observation in App. A.4, along with empirical results from Huh et al. (2024); Fan et al. (2025), suggests that increasing the capacity of unimodal models improves the alignment between unimodal representations, leading to better SUE performance without requiring additional paired data. Finally, this work establishes a foundation for a fully unpaired multimodal learning framework across arbitrary modalities, which we consider an intriguing and promising avenue for future work.
@@ Appendix A Additional Results
Our code, including SUE implementation and all experiments, will be publicly available upon acceptance.
@@
While the Flickr30k T2I task results are reported in the paper by Huang et al. (2022), the absence of public implementation prevents a more comprehensive comparison with MACK.
@@ A.1 Domain Confusion results
A.2 Zero-shot
To demonstrate SUE’s ability to capture high-level cross-modal semantics, we consider the zero-shot scenario Radford et al. (2021). This is a by-product of SUE, with no further adjustment. We use the zero-shot scenario to demonstrate the universality properties of SUE. Specifically, we train SUE on Flickr30k and use it to encode new ImageNet images Deng et al. (2009) and their labels. We then compute the cosine similarity between each image and all sentence labels, selecting the highest-scoring label. As shown in Fig. 6, SUE effectively labels images from the unobserved dataset, despite having only minimal image-text correspondence available during training. Tab. 3 further quantifies these results with a small 50-sample test set, compared to CLIP Radford et al. (2021), trained on 400 million pairs. (see App. D.3 for more details).
Accuracy
CLIP 99.26
SUE (ours) 88.31
A.3 Emergent Cross-domain Classification
Tab. 5 presents classification accuracy results (%) on the Office31 dataset Saenko et al. (2010), which contains 31 classes and three domains. We used the following two domains: Amazon images and DSLR images. In one experiment, we trained a classifier on the embeddings from the Amazon domain and used it to classify the DSLR domain, and vice versa. SUE achieves high performance with relatively few paired samples, outperforming the MMD baseline (see App. A.1 for more details on this baseline). For additional information on the dataset, its preparation, and classifier, please refer to App. D.4.
@@ A.3 Emergent Cross-domain Classification
#paired Flickr30k
I2T T2I
SUE (ours) 500 0.667 0.638
CLIP $\sim$ 400M 0.971 0.973
A.4 Effect of Different Unimodal Encoders
@@
A.5.1 Similarity of Random Walks on Corresponding Modalities (Fig. 1a)
@@
A.5.2 Universal Properties of Spectral Embedding (Fig. 1b)
In this experiment, we investigated whether the eigenfunctions of the Laplace-Beltrami operator exhibit universal properties across modalities. Specifically, we asked whether the eigenfunctions of independently constructed Laplace-Beltrami operators on each unimodal representation, $\psi$ and $\phi$, are similar. To test this, we computed the first three eigenfunctions for each modality in the MSCOCO dataset using SpectralNet Shaham et al. (2018), and aligned them using CCA on a paired set. Fig. 1b presents the cosine distances between pairs compared to the cosine distances between non-pairs. shows the cosine distances between aligned pairs compared to randomly matched non-pairs. Notably, paired samples are significantly closer in the aligned space, indicating that for a pair $(x,x^{\prime})$, $\psi(x)\sim\phi(x^{\prime})$.
A.6 Extended Comparison to Contrastive with Various #paired

### v2

@@ 5 Experiments
The remainder of this section is organized as follows. Sec. 5.1 evaluates retrieval performance. Sec. 5.2 presents a range of applications of universality. To better understand the contribution of each component in SUE, we investigate the effects of various components. In Sec. 5.3, we examine the impact of incorporating additional paired and unpaired samples, revealing that unpaired data holds greater potential than previously recognized. Sec. 5.4 presents an ablation study assessing the roles of SUE’s components (SE, CCA, and MMD), demonstrating that the novel integration of these techniques is central to SUE’s effectiveness. Due to space constraints, we refer to App. A for additional results. Further details on the experimental setup, implementation, and training procedures are provided in Appendix E.
@@
To evaluate SUE’s performance, we use several paired datasets. To provide informative qualitative results and facilitate an intuitive understanding of the universal embedding concept, we use three vision-language datasets ([Uncaptioned image]): Flickr30k [59], MSCOCO [50], and Polyvore [31]; as well as a vision-vision dataset ([Uncaptioned image]): Edges2Shoes [37]; and a tabular-to-tabular dataset ([Uncaptioned image]): Handwritten [20]. For the generation task, we use another vision-language dataset ([Uncaptioned image]): caption-FFHQ [40]. See App. E.1 for further details on the datasets and preprocessing (e.g., the unimodal models used and adaptation to weakly-paired settings).
@@
To the best of our knowledge, this is one of the very first works to exploit weakly-paired data (i.e., very few pairs). As such, no direct baselines exist for this setting. To contextualize results, we compare SUE with the contrastive method designed for fully paired data [60, 80, 22] (training an MLP on pre-trained unimodal features using the same number of pairs), and the CSA method [48] designed for small paired datasets. In addition, we examine domain confusion (App. A.1) and SAIL [83] (App. A.2), a representative approach for incorporating unpaired data alongside paired data; both perform notably worse in these scenarios.
Table 1: Retrieval results. Results with few paired samples on vision-language ([Uncaptioned image]), vision-vision ([Uncaptioned image]), and tabular-tabular ([Uncaptioned image]) datasets from each modality to another: image-to-text (I2T), text-to-image (T2I), edges-to-shoes (E2S), shoes-to-edges (S2E), KL coefficients-to-pixel averages (K2P), pixel-to-KL (P2K); by SUE, Contrastive, and CSA. The Imp. column states the relative mean improvement of SUE over Contrastive learning. Using the same small number of pairs, SUE significantly outperforms the popular paired method. SUE substantially relies on unpaired data.
#paired SUE (ours) Contrastive CSA Imp. R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 MSCOCO 100 I2T 5.75 21.50 34.25 1.50 8.50 13.00 0.00 1.25 3.00 +257.20% [Uncaptioned image] T2I 5.25 18.25 33.25 0.80 5.80 12.20 0.00 1.00 2.25 Flickr30k 500 I2T 4.25 19.75 32.00 3.00 9.50 16.20 0.25 1.25 2.50 +103.32% [Uncaptioned image] T2I 5.75 22.00 32.75 2.50 9.80 15.00 0.25 0.75 2.75 Polyvore 500 I2T 6.00 22.75 32.25 3.20 13.8 22.5 0.25 1.25 2.25 +55.67% [Uncaptioned image] T2I 4.75 20.75 32.00 4.00 11.50 23.00 0.25 1.00 3.25 Edges2Shoes 50 E2S 4.00 16.00 25.25 1.00 5.50 14.00 0.25 1.50 2.75 +200.51% [Uncaptioned image] S2E 3.50 17.00 27.00 0.80 6.00 12.80 0.25 1.50 3.00 Handwritten 100 K2P 25.50 62.00 79.00 4.80 17.00 28.00 4.25 12.25 17.50 +283.60% [Uncaptioned image] P2K 25.00 61.75 78.00 4.80 17.80 30.50 3.50 9.00 15.00
@@ 5.1 Retrieval
Tab. 1 reports retrieval scores for SUE and contrastive learning on several datasets using a randomly sampled test set of 400 examples. Following prior work [60, 22], we use the Recall@$k$ metric. The results demonstrate SUE’s ability to capture cross-modal semantics with as few as 100 pairs for MSCOCO, 500 for Flickr30k and Polyvore, and just 50 for Edges2Shoes. Remarkably, these results are also comparable with the ones reported by Huang et al. [35] on the Flickr30k T2I task [^1] R@1 10.3; R@5 25.1; R@10 34.0., despite MACK using a paired segmentation-to-text model (see App. A). In App. A.5 we report mean average precision (mAP) results, reinforcing the same conclusions. In App. A.6, we further analyze how different unimodal encoders affect SUE’s retrieval performance.
@@
In Fig. 4a we supplement the quantitative results with several qualitative examples. Additional qualitative results can be found in App. A.9. These results illustrate how SUE successfully captures semantic relationships across modalities, even in cases where the exact pair is not among the first neighbors (leading to a zero Recall@$k$ score).
In Fig. 2, we consider a custom caption scenario, where we retrieve images of a more general query sentence that cannot be found in the dataset. These examples further support SUE’s ability to capture these cross-modal semantics. These qualitative examples also depict the challenge of reliably measuring semantic coherence across modalities in retrieval tasks using the recall@$k$ metric. Therefore, in App. C we consider a new soft variant of Recall@$k$, which follows previous works [33] and uses CLIP score as a measure of semantic alignment between image-text pairs.
@@ 5.2 Applications of Universality
Fig. 4b depicts several examples of text-to-image generation with minimal text-image correspondence. In particular, we use a pre-trained GAN inversion model [77] trained on the FFHQ dataset [40] and train a converter model that maps from the universal image embeddings to the GAN’s latent space. During inference, we pass universal text embeddings to the converter and generate face images from its output using the GAN decoder. Noteworthy, both the GAN and the converter were trained exclusively on images, with few text-image pairs available only during SUE’s training. For more details on the training process of the converter model, please refer to App. E.3.
@@
Fig. 4c presents the results of emergent arithmetic operations on the universal embedding. Notably, these examples require no additional training - we simply perform vector summation of text and image embeddings and generate the image using the pre-trained converter and GAN. This approach aligns with techniques demonstrated using CLIP [61] but with limited text-image pairing. This summation in the vector space translates into an intuitive integration of semantics, making it, in a sense, a semantic vector space, while almost no text-image pairs were available during the construction of the shared embedding. This showcases the SUE’s cross-modality capturing abilities, even in the absence of large paired datasets.
@@
In App A.3 we further demonstrate SUE’s ability to capture high-level cross-modal, using a zero-shot scenario - acting as a classifier without being explicitly trained for classification.
@@
In App A.4 we demonstrate the effectiveness of universality for the classification of an unlabeled domain with minimal correspondence to a labeled domain.
@@ 5.4 Ablation Study
Tab. 2 presents the results of various ablations of SUE, from the original unimodal representations to the complete SUE pipeline. Alongside these empirical results, we provide UMAP visualizations of SUE’s three steps in Fig. 13 in App. A.11. The main outcome from these results is that SE has a pivotal role in revealing the universal embedding. Attempts to align the original unimodal representations yield mostly trivial retrieval results (see left column of each dataset in Tab. 2). However, applying the same alignment approach (i.e., CCA alongside MMD) on the SE space, significantly improves the retrieval performance. This underscores SE’s essential role in extracting the modality-invariant representations. This is further supported by the results in Tab. 10 in App. A.10, where replacing the SE step with an Auto-Encoder significantly decreases the performance. These results support the existence of the universal embedding. SUE successfully learns a meaningful cross-modal semantic representation with only a minimal number of paired samples, used solely to facilitate linear cross-modal alignment. See extended ablation study in App. A.10.
@@ 6 Conclusion and Future Work
In doing so, we challenge the prevailing approach in multimodal learning, which relies heavily on extremely large amounts of paired samples. SUE, operating with weakly-paired data, in the sense of very few pairs, yields meaningful results across various applications, especially highlighted by the comparison to contrastive learning.
Limitations.
Despite these contributions, several limitations remain. While SUE provides compelling evidence for the hidden potential of unpaired data in learning a shared representation space, it does not yet match the performance of state-of-the-art models trained with massive paired datasets. Furthermore, our evaluation has been limited to a small set of modalities and tasks, leaving open questions about its generalizability to more complex domains such as video, speech, or scientific data. Addressing these challenges offers a promising avenue for advancing the framework.
Importantly, our work does not yet aim to replace existing multimodal models trained with large sets of paired data. Rather, it introduces a new concept for settings where such supervision is unavailable. Interestingly, the observation in App. A.6, along with empirical results from Huh et al. [36], Fan et al. [25], suggests that increasing the capacity of unimodal models improves the alignment between unimodal representations, leading to better SUE performance without requiring additional paired data. Finally, this work establishes a foundation for a fully unpaired multimodal learning framework across arbitrary modalities, which we consider an intriguing and promising avenue for future work.
@@ Appendix A Additional Results
Our code, including SUE implementation and all experiments, is available at https://github.com/shaham-lab/SUE.
@@
While the Flickr30k T2I task results are reported in the paper by Huang et al. [35], the absence of public implementation prevents a more comprehensive comparison with MACK.
@@ A.1 Domain Confusion results
A.2 Comparison with SAIL
Tab. 4 presents a comparison between SUE and SAIL [83], a representative method for incorporating large unpaired data alongside paired data. The comparison is on vision-language datasets ([Uncaptioned image]) under the same settings as Tab. 1. The results demonstrate that SUE is more effective in the weakly paired scenario.
Table 4: SAIL comparison results. Results with few paired samples on vision-language ([Uncaptioned image]) from each modality to another: image-to-text (I2T), text-to-image (T2I); by SUE and SAIL. Using the same small number of pairs, SUE significantly outperforms SAIL.
#paired SUE (ours) SAIL
R@1 R@5 R@10 R@1 R@5 R@10
MSCOCO 100 I2T 5.75 21.50 34.25 0.75 4.75 7.25
[Uncaptioned image] T2I 5.25 18.25 33.25 0.75 4.50 7.00
Flickr30k 500 I2T 4.25 19.75 32.00 1.00 3.75 6.50
[Uncaptioned image] T2I 5.75 22.00 32.75 0.50 3.00 6.75
Polyvore 500 I2T 6.00 22.75 32.25 0.50 4.25 7.25
[Uncaptioned image] T2I 4.75 20.75 32.00 0.75 4.00 7.50
A.3 Zero-shot
To demonstrate SUE’s ability to capture high-level cross-modal semantics, we consider the zero-shot scenario [60]. This is a by-product of SUE, with no further adjustment. We use the zero-shot scenario to demonstrate the universality properties of SUE. Specifically, we train SUE on Flickr30k and use it to encode new ImageNet images [17] and their labels. We then compute the cosine similarity between each image and all sentence labels, selecting the highest-scoring label. As shown in Fig. 6, SUE effectively labels images from the unobserved dataset, despite having only minimal image-text correspondence available during training. Tab. 5 further quantifies these results with a small 50-sample test set, compared to CLIP [60], trained on 400 million pairs. (see App. E.3 for more details).
Accuracy
CLIP 99.26
SUE (ours) 88.31
A.4 Emergent Cross-domain Classification
Tab. 6 presents classification accuracy results (%) on the Office31 dataset [64], which contains 31 classes and three domains. We used the following two domains: Amazon images and DSLR images. In one experiment, we trained a classifier on the embeddings from the Amazon domain and used it to classify the DSLR domain, and vice versa. SUE achieves high performance with relatively few paired samples, outperforming the MMD baseline (see App. A.1 for more details on this baseline). For additional information on the dataset, its preparation, and classifier, please refer to App. E.4.
@@ A.4 Emergent Cross-domain Classification
A.5 mAP Results
In Tab. 7 we provide the mean Average Precision (mAP) results, in the same settings as in Tab. 1. SUE outperforms both contrastive and CSA under this ranking-sensitive metric as well, further validating its effectiveness.
Table 7: mAP results. Results with few paired samples on vision-language ([Uncaptioned image]), vision-vision ([Uncaptioned image]), and tabular-tabular ([Uncaptioned image]) datasets from each modality to another: image-to-text (I2T), text-to-image (T2I), edges-to-shoes (E2S), shoes-to-edges (S2E), KL coefficients-to-pixel averages (K2P), pixel-to-KL (P2K); by SUE, Contrastive, and CSA. Using the same small number of pairs, SUE significantly outperforms the popular paired method.
#paired SUE (ours) Contrastive CSA
mAP mAP mAP
MSCOCO 100 I2T 9.72 5.00 1.48
[Uncaptioned image] T2I 9.80 4.90 1.38
Flickr30k 500 I2T 8.91 5.80 1.67
[Uncaptioned image] T2I 8.07 5.80 1.67
Polyvore 500 I2T 9.87 8.8 1.54
[Uncaptioned image] T2I 9.35 9.2 1.68
Edges2Shoes 50 E2S 7.03 5.7 1.77
[Uncaptioned image] S2E 6.99 5.1 1.88
Handwritten 100 K2P 14.38 12.60 9.42
[Uncaptioned image] P2K 14.03 12.90 7.96
A.6 Effect of Different Unimodal Encoders
@@
To account for recent advances in text encoders, we additionally evaluate GTR [57], a larger and stronger text encoder, in Tab. 8. Interestingly, its absolute performance is slightly lower. We hypothesize that the Mini-LM was already sufficient given the strength of the vision encoder (DINOv2), or alternatively, that GTR is less aligned with the image representations. In either case, the core effectiveness of our method appears robust to the choice of text encoder.
@@
A.7.1 Similarity of Random Walks on Corresponding Modalities (Fig. 1a)
@@
A.7.2 Universal Properties of Spectral Embedding (Fig. 1b)
In this experiment, we investigated whether the eigenfunctions of the Laplace-Beltrami operator exhibit universal properties across modalities. Specifically, we asked whether the eigenfunctions of independently constructed Laplace-Beltrami operators on each unimodal representation, $\psi$ and $\phi$, are similar. To test this, we computed the first three eigenfunctions for each modality in the MSCOCO dataset using SpectralNet [67], and aligned them using CCA on a paired set. Fig. 1b presents the cosine distances between pairs compared to the cosine distances between non-pairs. shows the cosine distances between aligned pairs compared to randomly matched non-pairs. Notably, paired samples are significantly closer in the aligned space, indicating that for a pair $(x,x^{\prime})$, $\psi(x)\sim\phi(x^{\prime})$.
A.8 Extended Comparison to Contrastive with Various #paired